OpenStudy (anonymous):

PLEASE HALP Write an equation of the line, in point-slope , that passes through the two given points. points: (-2, 15), (9, -18)

OpenStudy (anonymous):

Hello?

OpenStudy (tkhunny):

You'll need the slope between those two points. Please demonstrate.

OpenStudy (***[isuru]***):

hi, to find the equation u must firs find out the gradient and the intercept... If we assume that the format of the equation we will get is y = mx + c where "m" is the gradient or the slope and "c" is the intercept... Let's find the slop first... \[gradient = \frac{ \Delta Y }{ \Delta X }\] where.. \[\Delta Y = difference \ of \ Y \ coordinates\] \[\Delta X = difference \ of \ x \ coordinates\] so with the 2 pair of coordinates given the difference of X coordinates is = 9 - (-2) = 11 difference of Y coordinates is = -18 - (15) = -33 so ... gradient = -33/11 = -3 so we can rewrite our equation as... y = mx + c y =-3)x + c now to find the intercept let's use the coordinate (-2, 15), according to that y coordinate = 15 x coordinate = -2 let's substitute them in our equation... y = -3x +c 15 = -3(-2) +c c = 9 that's it!! we have now finished the equation of the question.... So the finale answer is y = -3x +9 Hope this will help ya!!!

OpenStudy (***[isuru]***):

Did ya get it :-) ?